In class, we said that a 1000-500 hPa thickness of 5400m represents a 50% probability of rain versus snow. What mean column temperature does this represent in degrees Celsius? (Use a value of 9.8 for gravity)
Answer #1

from hydrostatic equation

Z2-Z1=RT/g*ln(p1/p2)

where,

R= gas constant=287J/Kg/K

T =Mean temperature

Z1,Z2 (m)= height of isobaric surfaces p1,p2

g= 9.81m/s^2

given in question

p1= 1000hPa, p2=500hPa , Z2-Z1= 5400m

SO,

5400m= 287J/Kg/K*T/9.81m/s^2* ln(1000hPa/500hPa)

T= 266.29 C